Beginning in 2011, city hall, administrative offices, and municipal courts in city of EI Paso, Texas, will go on a 10 hour/day, 4-day workweek from the beginning of May through the end of September. The shortened workweek will affect 25% of the city’s 6100 employees and will save $42,600 per month for those 5 months, because of reduced fuel, utility, and janitorial expenses. If this work schedule continues for next 10 years, what is the future worth of the savings at the end of that time (i.e., end of year 2020)? Use an interest rate of 0.5% per month.
Answer #1

Monthly saving = $42,600

Number of months = 12 x 10 = 120

Future worth ($) = 42,600 x F/A(0.5%, 120) = 42,600 x 163.8793** = 6,981,258.18

**F/A(0.5%, 120) = [(1.005)120 - 1] / 0.005 = (1.8194 - 1) / 0.005 = 0.8194 / 0.005 = 163.8793